Owl Shift, 6/6/99

Peter Shanahan (shanahan@fnal.gov)
Sun, 6 Jun 1999 06:55:36 -0500

Shift Summary for Owl shift, 06/Jun/99

On Shift:

Hours of beam delivered at all intensities: 0
Hours of beam delivered at nominal intensity: 0
Hours of E832 data (special runs not included): 0
Hours of special runs (laser scans, muon runs...): 7

Accounting of lost time: E832 Switchover.

========================================================================
Runs ended this shift:
Run 13388 - 11M events - COSMIC (to disk only)

Tapes brought to FCC this shift: None
(including tapes for online split)

========================================================================
Accesses during shift: Open access. Hogan arrives 0600 to prepare
for DC survey.

========================================================================
Description of Shift:

run spill time comment
------------------------------------------------------------------------
13388 0135 Ordinal number mismatch, crate 65.
13388 0239 Gateway hangs, PTI VME bus error.
Reset plane 3 VME crate, fresh start.
--> DDD init failed. Reset/fresh start, finally OK.
13392 0434 Unknown message received from gateway, but
run continues OK.